### v1

# Neutrinoless Double Beta Decay in Light of JUNO First Data

Abstract
The first results from the JUNO reactor neutrino oscillation experiment improve our knowledge of neutrino masses and mixing parameters, especially the solar angle $\theta_{s}\equiv\theta_{12}$ and the solar mass squared difference $\Delta m^{2}_{s}\equiv\Delta m^{2}_{21}$. We discuss the implications of these results on neutrinoless double beta decay by itself and in combination with the global fit of neutrino oscillation experiments, the JUNO first data, and cosmological constraints on the neutrino mass sum. For the effective mass $\langle m_{ee}\rangle$, the uncertainties in its lower limits for both mass orderings and upper limits for the normal ordering are largely reduced. Since the cosmological CMB and DESI BAO data put a stringent constraint on the neutrino mass scale, we also show how the probability distribution of both the real and imaginary parts of the effective mass $\langle m_{ee}\rangle$ on the complex plane is affected. Especially, the funnel region with $|\langle m_{ee}\rangle|\lesssim 1$ meV receives larger chance to happen. Correspondingly, the chance of determining the two Majorana CP phases simultaneously in this region also increases with reduced uncertainty.
1 Introduction
The baryon asymmetry of the Universe, namely that almost no anti-matter exists in Nature [Cohen98, Canetti12], is one of the main problems in contemporary astroparticle physics and cosmology. An inflationary phase at the beginning of our Universe dilutes all pre-existing matter [Guth81] which requires a mechanism to understand the observed baryon asymmetry. The Sakharov conditions [Sakharov:1967dj] summarize the essential ingredients which must be fulfilled. One very well motivated mechanism is leptogenesis [Fukugita86] which is typically related to Majorana neutrino masses with additional CP phases emerging by the seesaw mechanism [Fritzsch:1975sr, Minkowski:1977sc, Yanagida:1979as, Gell-Mann:1979vob, Glashow:1979nm, Mohapatra:1979ia, Konetschny:1977bn, Magg:1980ut, Schechter:1981cv, Cheng:1980qt, Lazarides:1980nt, Mohapatra:1980yp, Foot:1988aq, Ma:1998dn].
Verifying the Majorana nature of neutrinos is therefore very important for justifying the vanilla leptogenesis mechanism. Although the ultra-violet physics at high energy is difficult to observe with current technology, with two exceptions of cosmic string [Dror:2019syi] and cosmological collider [Cui:2021iie], the nature inherited by light active neutrinos can manifest itself in the neutrinoless double beta decay ($0\nu 2\beta$) process where a mother nuclei decays into a daughter nuclei plus two electrons, $^{Z}_{A}N\rightarrow{}^{Z+2}_{A}N+2e^{-}$ [Dolinski:2019nrj, Shergold:2021evs, Agostini:2022zub, Cirigliano:2022oqy]. With two electrons appearing in the final state, such process indicates that the lepton number is violated by two. In the minimal scenario, the $0\nu 2\beta$ decay can be induced by an effective Majorana mass $m_{ee}$ of the light active neutrinos in the electron flavor. Reciprocally, a lepton number violating process would finally contribute to a light Majorana mass [Schechter82].
However, the $0\nu 2\beta$ decay process involves not just particle physics, but also nuclear reaction and atomic effects. The decay half-life can be expressed as
$\left(T^{0\nu}_{1/2}\right)^{-1}\equiv G^{0\nu}|M^{0\nu}|^{2}|\langle m_{ee}\rangle|^{2},$ (1.1)
where $G^{0\nu}$ is the phase space factor [Kotila:2012zza, Kotila:2013gea, Mirea:2014dza, Neacsu:2015uja] associated with the decay kinematics and atomic physics, $M^{0\nu}$ is the nuclear matrix element (NME), and $\langle m_{ee}\rangle$ is the effective electron neutrino mass from particle physics. Of these three parts, the phase space factor can be calculated very precisely while the NME suffers a lot from theoretical uncertainties [Agostini:2022zub]. The NME calculations involve complex nuclear many-body physics and depend on the nuclear structure models. There are multiple theoretical approaches such as the quasiparticle random phase approximation (QRPA) [qrpa1, qrpa2, qrpa3, qrpa4, qrpa5], interacting shell mode (ISM) [shell1, shell2, shell3, shell4], energy density functional methods (EDF) [edf1, edf2, edf3], interacting boson model (IBM) [ibm1, ibm2], and ab initio techniques [abinitio1, abinitio2, abinitio3, abinitio4, abinitio5, abinitio6, abinitio7]. Their predictions typically differ by a factor of $2\sim 3$, representing the main source of systematic uncertainty from the nuclear physics side [Agostini:2022zub].
On the other hand, the particle physics side has improved quite significantly. The improvement comes from neutrino oscillation experiments, beta decay spectrum measurement around the end point, and cosmology. Especially, there is complementarity between the reactor neutrino oscillation experiments and $0\nu 2\beta$ decay [Pascoli:2005zb, Choubey:2005rq, Lindner:2005kr, Dueck:2011hu, GE1, GE2, Ge:2019ldu] since both involves only the electron flavor and share the same set of neutrino oscillation parameters. To be exact, only two mixing angles, the reactor angle $\theta_{r}\equiv\theta_{13}$ and the solar angle $\theta_{a}\equiv\theta_{12}$ can appear in the effective Majorana mass $m_{ee}$ for $0\nu 2\beta$ decay and the disappearence probability $P_{ee}$ of reactor neutrino oscillation. In addition, the solar mass squared difference $\Delta m^{2}_{s}\equiv\Delta m^{2}_{21}$ and the atmospheric one $\Delta m^{2}_{a}\equiv\Delta m^{2}_{31}$ can also enter. It is interesting to see that these four oscillation parameters can be precisely measured with the combination of Daya Bay and JUNO [Ge:2012wj, JUNO:2015zny]. Especially, the recent global fits [NuFIT6] and JUNO first data [JUNO1st] have significantly reduced the uncertainty of the solar angle $\theta_{s}$. In addition, the KATRIN experiment provides the best measurement of the beta decay spectrum to put a $\sum_{i}m_{i}\leq 0.45$ eV limit at 90% C.L. (confidence level) [KATRIN2022, KATRIN2024]. The most stringent bound on the absolute neutrino mass comes from the recent DESI BAO data [DESI2024, DESI2024b], plus other existing observations such as CMB [Planck:2018vyg].
Concerning the recent progress in neutrino experiments and cosmology, we would like to show how these affect the prospects of the $0\nu 2\beta$ decay. We first summarize the particle physics contribution to the $0\nu 2\beta$ decay in Sec. 2. The following Sec. 3 studies the influence of the recent global fit and JUNO first data release in particular on the effective mass and half-life time. We further study the possibility for the effective mass $\langle m_{ee}\rangle$ to fall into the funnel region and the prospect of determining the two Majorana CP phases simultaneously in Sec. 4. The conclusion of this paper can be found in Sec. 5.
2 Effective Majorana Mass for $0\nu 2\beta$ Decay
As summarized in Eq. (1.1), the half-life time $T^{0\nu}_{1/2}$ contains three parts from atomic physics, particle physics, and nuclear physics, respectively. This section summarizes the effective mass $\langle m_{ee}\rangle$ from the particle physics side. With $\mathcal{O}(1)$ MeV energy release, the $0\nu 2\beta$ process can only produce two electrons which explain why only $\langle m_{ee}\rangle$ for the electron flavor can be probed.
The effective mass $\langle m_{ee}\rangle$ is defined as:
$\langle m_{ee}\rangle\equiv\sum_{i}m_{i}U_{ei}^{2},$ (2.1)
where $m_{i}$ are the neutrino mass eigenvalues and $U_{ei}$ are the PMNS mixing matrix elements. We follow the PMNS matrix parametrization [PDG24],
[…]
with $\mathcal{P}\equiv\text{diag}\{e^{-i\beta_{1}},e^{-i\beta_{2}},e^{-i\beta_{3}}\}$ and $\mathcal{Q}\equiv\text{diag}\{e^{-i\delta_{\rm M1}/2},e^{-i\delta_{\rm M2}/2},e^{-i(\delta_{\rm M3}-\delta_{D})/2}\}$. The three phases $\beta_{i}$ in $\mathcal{P}$ are unphysical while $\mathcal{Q}$ contains two physical Majorana phases. We set $\delta_{\rm M2}=0$ but take $\delta_{\rm M1}$ and $\delta_{\rm M3}$ as the two independent physical Majorana phases. The benefit of using such parametrization is that when the lightest mass vanishes, $m_{1}=0$ for the normal ordering (NO) case or $m_{3}=0$ for the inverted ordering (IO), one physical Majorana CP phase would automatically disappear altogether. For convenience, we labeled the three mixing angles according to the major type of experiments that fixes their values, $\theta_{a}\equiv\theta_{23}$ for the atmospheric angle, $\theta_{r}\equiv\theta_{13}$ for the reactor angle, and $\theta_{s}\equiv\theta_{12}$ for the solar angle.
Note that the third element of $\mathcal{Q}$ is defined as $e^{-i(\delta_{\rm M3}-\delta_{D})/2}$ where the Dirac CP phase $\delta_{D}$ is to cancel the same Dirac CP phase associated with the reactor angle $s_{r}$ of $U$ as defined in Eq. (2.2),
$\langle m_{ee}\rangle=m_{1}|U_{e1}|^{2}e^{i\delta_{\rm M1}}+m_{2}|U_{e2}|^{2}+m_{3}|U_{e3}|^{2}e^{i\delta_{\rm M3}}.$ (2.3)
A particularly illuminating perspective emerges when treating $\langle m_{ee}\rangle$ as a vector sum in the complex plane which can give a geometrical view [Vissani:1999tu, Xing:2014yka, GE2]. We can decompose it into three contributions,
$\langle m_{ee}\rangle\equiv\overrightarrow{L_{1}}+\overrightarrow{L_{2}}+\overrightarrow{L_{3}},$ (2.4)
with each vector being defined as,
$\overrightarrow{L_{1}}\equiv m_{1}|U_{e1}|^{2}e^{i\delta_{\rm M1}},\qquad\overrightarrow{L_{2}}\equiv m_{2}|U_{e2}|^{2},\qquad\overrightarrow{L_{3}}\equiv m_{3}|U_{e3}|^{2}e^{i\delta_{\rm M3}}.$ (2.5)
With our phase convention $\delta_{\rm M2}=0$, $\overrightarrow{L_{2}}$ lies along the real axis while $\overrightarrow{L_{1}}$ and $\overrightarrow{L_{3}}$ rotate with the corresponding Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$, respectively. This geometric representation will be discussed in detail in Sec. 4, where we explore how the interplay between these vectors constrains the allowed values of $\langle m_{ee}\rangle$.
Using the specific parametrization of the PMNS matrix as defined in Eq. (2.2), we can write the mixing matrix elements explicitly,
$\langle m_{ee}\rangle=c^{2}_{s}c^{2}_{r}m_{1}e^{i\delta_{\rm M1}}+s^{2}_{s}c^{2}_{r}m_{2}+s^{2}_{r}m_{3}e^{i\delta_{\rm M3}},$ (2.6)
where $(c_{\alpha},s_{\alpha})\equiv(\cos\theta_{\alpha},\sin\theta_{\alpha})$. Note that the atmospheric mixing angle $\theta_{a}$ does not appear explicitly in $\langle m_{ee}\rangle$, which avoids any ambiguity related to the determination of the octant of $\theta_{a}$. The $0\nu 2\beta$ decay shares four oscillation parameters with the reactor neutrino oscillation. Besides the reactor angle $\theta_{r}$ and the solar angle $\theta_{s}$, both the solar mass squared difference $\Delta m^{2}_{s}\equiv m^{2}_{2}-m^{2}_{1}$ and the atmospheric mass squared difference $\Delta m^{2}_{a}\equiv|\Delta m^{2}_{3\ell}|=|m^{2}_{3}-m^{2}_{\ell}|$ where $\ell=1$ for NO ($m_{1}<m_{2}<m_{3}$) and $\ell=2$ for IO ($m_{3}<m_{1}<m_{2}$) also appear. To make the physics behind the mass squared differences more explicit, we have also adopted their labels according to the major type of observations from which their values are determined.
The behavior of $\langle m_{ee}\rangle$ significantly depends on the ordering of the neutrino mass, with fundamentally different phenomenological implications for each scenario. The main distinction arises from the lightest mass, $m_{1}$ for NO and $m_{3}$ for IO. This difference has direct consequences for the allowed range of $\langle m_{ee}\rangle$, particularly in the limit where the lightest mass is smaller than $0.1$ eV. Note that distinguishing the mass ordering is the major physics goal of the JUNO experiment [JUNO:2015zny].
For NO, we express the heavier masses ($m_{2}$ and $m_{3}$) in terms of $m_{1}$ and the measured mass-squared differences,
$\langle m_{ee}\rangle=m_{1}c^{2}_{s}c^{2}_{r}e^{i\delta_{\rm M1}}+\sqrt{\Delta m^{2}_{s}+m_{1}^{2}}c^{2}_{r}s^{2}_{s}+\sqrt{\Delta m^{2}_{a}+m_{1}^{2}}s^{2}_{r}e^{i\delta_{\rm M3}}.$ (2.7)
In this case, $\langle m_{ee}\rangle$ can vanish if the Majorana phases conspire to produce a complete destructive interference between the remaining contributions. Thus, observing $\langle m_{ee}\rangle=0$ would not exclude the Majorana nature of neutrinos under NO.
For IO, the masses are parametrized in terms of the lightest mass $m_{3}$:
$\langle m_{ee}\rangle=\sqrt{m_{3}^{2}+\Delta m^{2}_{a}-\Delta m^{2}_{s}}c^{2}_{s}c^{2}_{r}e^{i\delta_{\rm M1}}+\sqrt{m_{3}^{2}+\Delta m^{2}_{a}}c^{2}_{r}s^{2}_{s}+m_{3}s^{2}_{r}e^{i\delta_{\rm M3}}.$ (2.8)
The main feature of IO is that $\langle m_{ee}\rangle$ possesses a non-zero lower bound. This lower bound, which depends on the mixing angles and mass splitting, provides a testable prediction that can potentially discriminate between mass orderings for Majorana neutrinos. The minimum values allowed for $\langle m_{ee}\rangle$ for IO will be explored in the next Sec. 3.
[…]
The interplay between the neutrino oscillation parameters and the predictions for neutrinoless double-beta decay has been extensively studied in the literature, with detailed analysis of how experimental uncertainties propagate into the allowed ranges for the effective Majorana mass $\langle m_{ee}\rangle$ [Pascoli:2002xq, Pascoli:2005zb, Choubey:2005rq, Lindner:2005kr, Dueck:2011hu, GE1]. In this section, we perform an updated analysis of [GE1] by incorporating the latest global fit results from neutrino oscillation experiments [NuFIT6] and more crucially the first measurement of solar oscillation parameters $\sin^{2}\theta_{s}$ and $\Delta m^{2}_{s}$ from the JUNO experiment [JUNO1st]. Our goal is to quantify how the improved determination of the solar mixing angle $\theta_{s}$ from the JUNO initial 59.1-day dataset impacts the theoretical predictions of $\langle m_{ee}\rangle$.
As commented in the previous Sec. 2, different neutrino mass orderings have very different phenomenological consequences on the effective $\langle m_{ee}\rangle$ of $0\nu 2\beta$ decay. While $\langle m_{ee}\rangle$ could vanish for NO, it is bounded from not just above but also below for IO,
$\langle m_{ee}\rangle^{\rm IO}_{\rm max}$ $\equiv\sqrt{m^{2}_{3}+\Delta m^{2}_{a}}c^{2}_{s}c^{2}_{r}+\sqrt{m^{2}_{3}+\Delta m^{2}_{a}-\Delta m^{2}_{s}}s^{2}_{s}c^{2}_{r}+m_{3}s^{2}_{r}\,$, (3.1a)
$\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ $\equiv\sqrt{m^{2}_{3}+\Delta m^{2}_{a}}c^{2}_{s}c^{2}_{r}-\sqrt{m^{2}_{3}+\Delta m^{2}_{a}-\Delta m^{2}_{s}}s^{2}_{s}c^{2}_{r}-m_{3}s^{2}_{r}\,.$ (3.1b)
The upper and lower limits happen when the two Majorana phases take specific values to give constructive or destructive interference, respectively.
As shown in the left panel of Fig. 1, these two limits become flat when the lightest mass, $m_{1}$ for NO and $m_{3}$ for IO, approaches zero. More generally, a vanishing lightest mass would not have much effect on the $0\nu 2\beta$ effective mass $\langle m_{ee}\rangle$. Then, only the two mixing angles ($\theta_{r}$ and $\theta_{a}$) and the two mass squared differences ($\Delta m^{2}_{a}$ and $\Delta m^{2}_{s}$) enters. For IO,
$\langle m_{ee}\rangle^{\rm IO}_{\rm max}\approx\sqrt{\Delta m^{2}_{a}}c_{r}^{2},\qquad\langle m_{ee}\rangle^{\rm IO}_{\rm min}\approx\sqrt{\Delta m^{2}_{a}}(c_{s}^{2}-s_{s}^{2})c_{r}^{2}.$ (3.2)
Although both mass squared differences are involved in the electron flavor effective mass $\langle m_{ee}\rangle$, the limits are dominated by the atmohsperic one $\Delta m^{2}_{a}$ with much larger value. Further, the upper limit $\langle m_{ee}\rangle^{\rm IO}_{\rm max}$ even becomes independent of the solar angle $\theta_{s}$. Since both the atmospheric mass squared difference $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ have been measured quite precisely with just roughly 1% error [NuFIT6], there is almost no uncertainty in the upper limit for IO. This has been clearly shown in Fig. 1 as the filled regions for IO have roughly the same upper boundary.
Fig. 1: The particle physics uncertainties for the $0\nu 2\beta$ decay effective mass $\langle m_{ee}\rangle$. For comparison and easy viewing, three panels are shown: (Left) overall features for both NO and IO; (Middle) zoomed-in linear plot for IO; (Right) zoomed-in linear plot of the low-mass horizontal branch for NO. The allowed regions of the effective mass $\langle m_{ee}\rangle$ have been demonstrated with four different uncertainty configurations: (1) the current $3\,\sigma$ range from NuFIT 6.0 global fit (orange); (2) the updated $3\,\sigma$ range with the JUNO first data release (blue); (3) the expected $3\,\sigma$ range with full JUNO (green); (4) the ultimate scenario with negligible uncertainty (red). The current best limit on the effective mass $\langle m_{ee}\rangle$ from KamLAND-Zen with 800 kg Xenon mass has been added as hashed region with solid boundaries while the expected sensitivities from CUPID, LEGEND1000 and nEXO are shown in the left plot as dashed line without uncertainties of nuclear matrix elements. The middel plot includes the sensitivity of LEGEND1000 including uncertainties of nuclear matrix elements.
For comparison, the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ depends on not just $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ but also the solar angle $\theta_{s}$. Although $\theta_{s}$ has also been measured very precisely with only 2% uncertainty [NuFIT6], the combination $\cos 2\theta_{s}$ that appears in $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ can vary by 1/3 within the $3\,\sigma$ range. Even in the left panel of Fig. 1 with logarithmic scale, the effect of improving the solar angle measurement with JUNO can be clearly seen. Since the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ corresponds to the maximal value of half-life time $T^{0\nu}_{1/2}$ according to Eq. (1.1), such uncertainty would have significant consequences when interpreting the experimental data. Any improvement in the determination of $\theta_{s}$ would significantly reduce the uncertainty on the lower bound of $\langle m_{ee}\rangle$ for IO [Pascoli:2002xq, GE1].
For NO, the effective mass in Eq. (2.7) is also bound from below and above with a vanishing lightest mass $m_{1}\rightarrow 0$,
$\langle m_{ee}\rangle^{\rm NO}_{\rm min}\approx\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}-\sqrt{\Delta m^{2}_{a}}s^{2}_{r},\qquad\langle m_{ee}\rangle^{\rm NO}_{\rm max}\approx\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}+\sqrt{\Delta m^{2}_{a}}s^{2}_{r}.$ (3.3)
These two limits contain two mass squared differences and two mixing angles. Since the reactor angle $\theta_{r}$ is not large, its consine is roughly $c_{r}\approx 1$. The relative size between the two terms in Eq. (3.3) is then determined by comparing the ratios $\Delta m^{2}_{s}/\Delta m^{2}_{a}$ and $s^{2}_{r}/s^{2}_{s}$. According to the global fit [NuFIT6], $\sqrt{\Delta m^{2}_{s}/\Delta m^{2}_{a}}\approx 0.17$ and $s^{2}_{r}/s^{2}_{s}\approx 0.07$ which means the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ dominates. Note that the solar angle $s^{2}_{s}$ appears in the first term. The current $3\,\sigma$ variation of $s^{2}_{s}$ can be as large as 23% [NuFIT6]. Although the solar mass squared difference $\Delta m^{2}_{s}$ itself varies by $\sim 15\%$ at $3\,\sigma$, its contribution to $\langle m_{ee}\rangle$ is only $\sim 7.5\%$. So $\theta_{s}$ is the dominant source of uncertainty for NO. With all factors combined, the total effect of those uncertainties from the current data can be clearly seen in the orange bands of the left and right panels of Fig. 1. The difference is much more visible for the lower limits. This is because the variation in the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ actually affects the central values as well and the whole NO branch in the limit of vanishing lightest mass moves vertically. While the lower limit $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ is different between the two terms in Eq. (3.3), the upper one $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$ is a sum with larger value. Then the relative variation is much smaller for the upper limit.
To make the influence of $\theta_{s}$ on the effective mass $\langle m_{ee}\rangle$ more visible, we have included two linear scale plots for IO and NO as the middle and right panels of Fig. 1. From this figure, it is also possible to infer from the orange bands both the minimum and maximum values of the effective mass for each mass ordering in the vanishing lightest mass scenario. Explicitly, their $3\,\sigma$ intervals are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(15.72\sim 22.28)\,\text{meV}$, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.188\sim 1.828)\,\text{meV}$, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.404\sim 4.028)\,\text{meV}$, resulting from the propagation of the oscillation parameter uncertainties. While these uncertainties appear modest in $\langle m_{ee}\rangle$, their influence on the $0\nu 2\beta$ experiments is much more significant [GE1]. Especially, the half-life time scales inversely with the effective mass squared, $T^{0\nu}_{1/2}\propto 1/|\langle m_{ee}\rangle|^{2}$ according to Eq. (1.1). The variation in $\langle m_{ee}\rangle$ would get amplified when propagating to the half-life time. Further, to reach the required sensitivity on $T^{0\nu}_{1/2}$ the expected experimental exposure scales [Avignone:2005cs], $M\times t\propto T^{0\nu}_{1/2}$ where $M$ is the detector target mass and $t$ is the running time, with the half-life time squared. Finally, the experimental exposure can differ by a factor of 4.3 for IO and 2.5 for NO. The improvement on the uncertainty from the particle physics side by JUNO is expected to significantly reduce the uncertainty in the experimental design and data interpretation.
Following the first data release from JUNO, which reported 59.1 days of data taking [JUNO1st], the precision on the solar angle $\theta_{s}$ and the solar mass splitting $\Delta m^{2}_{s}$ has been significantly improved to $\sin^{2}\theta_{s}=0.3092\pm 0.0087$ and $\Delta m^{2}_{s}=(7.50\pm 0.12)\times 10^{-5}\,$eV2. This corresponds to a reduction of uncertainty by approximately a factor of $0.741$ at $3\,\sigma$ than the global fit results from NuFIT 6.0 [NuFIT6]. According to Eq. (3.2), this translates to around 22% reduction in the uncertainty of $\langle m_{ee}\rangle^{\rm IO}_{\rm min}\propto c^{2}_{s}-s^{2}_{s}$ at $3\,\sigma$. This improvement was achieved through JUNO’s unique capability to observe the low-frequency mode of the reactor antineutrino oscillation [Ge:2012wj],
$P_{ee}\approx 1-c^{4}_{r}(c^{2}_{s}-s^{2}_{s})\sin^{2}\left(\frac{\Delta m^{2}_{s}L}{4E_{\nu}}\right).$ (3.4)
For simplicity, we have omitted those high-frequency mode modulated by $\Delta m^{2}_{a}$ which are important for the mass ordering measurement. The amplitude of the low-frequency mode is approximately $c^{2}_{s}-s^{2}_{s}$ which is exactly the one that appears in the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ for IO. With large statistics at JUNO, the solar angle can be measured with very high precision. At the same time, the mode frequency or equivalently the solar mass squared difference $\Delta m^{2}_{s}$ can also be measured with comparable precision. This impressive result within just 59.1 days of data taking demonstrates JUNO’s potential to achieve sub-percent precision on solar parameters with the full dataset, making it the leading experiment for solar mixing angle determination in the coming years.
The improved determination of $\theta_{s}$ and $\Delta m^{2}_{s}$ directly translates into enhanced constraints on the effective Majorana mass $\langle m_{ee}\rangle$, as shown in Fig. 1. It narrows the allowed band for $\langle m_{ee}\rangle$ in both mass orderings. Particularly, the lower bound $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ is most relevant for the sensitivity goals of next-generation $0\nu 2\beta$ experiments.
For comparison, Fig. 1 shows four uncertainty scenarios. First, the theoretical prediction with perfect knowledge of oscillation parameters (red) that represents the ultimate goal of mitigating the uncertainty from the particle physics side to a negligible extend [GE1]. Second, the current $3\,\sigma$ range with the pre-JUNO global fits (orange) that shows what other experiments have achieved altogether [NuFIT6]. Between these two cases, there is a huge space which is exactly the existing uncertainty from particle physics. In the middle, we have show two uncertainty scenario from the JUNO experiment. The blue band shows the reduced uncertainty with the JUNO first data release [JUNO1st]. The comparison between the orange (pre-JUNO) and blue (post-JUNO) bands clearly illustrates the impact of the improved solar angle measurement by JUNO with a noticeable reduction in the width of the allowed regions for both mass orderings. The green one corresponds to the expected sensitivity with 6 years of data collection [JUNO:2015zny].
From Fig. 1, it is possible to observe that the JUNO determination of $\theta_{s}$ sharpens the lower edge of the IO band (compare blue vs. orange), with total uncertainty in $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ at the $3\sigma$ C.L. by:
$\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(16.36\sim 21.48)\,\mathrm{meV}.$ (3.5)
For comparison, we have also shown those expected sensitivities from $0\nu 2\beta$ decay experiments as horizontal lines. The most stringent current bounds come from KamLAND-Zen (136Xe), yielding $T_{1/2}^{0\nu}>2.3\times 10^{26}$ yr at $90\%$ C.L., which translates to $\langle m_{ee}\rangle<(36\sim 156)$ meV depending on the nuclear matrix elements (NME) [KamLAND-Zen:2024eml, KamLAND-Zen:2022tow]. This bound is indicated as the hashed region with solid boundaries in the left and middle panels of Fig. 1. The current experiments are beginning to probe the IO parameter space, particularly in the region where $m_{\text{lightest}}\gtrsim 10$ meV. The next-generation $0\nu\beta\beta$ experiments, including nEXO [nEXO1, nEXO2], LEGEND-1000 [LEGEND, LEGEND:2021bnm], and CUPID [CUPID1, CUPID2], are projected to reach sensitivities of $\langle m_{ee}\rangle<(5\sim 12)$ meV, which would fully cover the IO prediction band and begin to probe the NO region.
Fig. 2: One-dimensional $\Delta\chi^{2}$ for the $0\nu 2\beta$ effective mas limits $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$ (left), $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$ (center), and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ (right) according to the neutrino oscillation global fit (blue) [NuFIT6] or the combined results with both global fit and JUNO (red) [JUNO1st].
To make the influence of the JUNO first data [JUNO1st] on the $0\nu 2\beta$ effective mass limits more explicit, Fig. 2 shows the one-dimensional $\Delta\chi^{2}$ profiles for $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$, $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$, and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ with vanishing lightest mass ($m_{\text{lightest}}\to 0$). In this limit, the residual uncertainty is contributed by the oscillation parameters $\Delta m^{2}_{a}$, $\Delta m^{2}_{s}$, $\theta_{s}$, and $\theta_{r}$, as discussed before. The gray dashed horizontal lines indicate the $1\,\sigma$, $2\,\sigma$, and $3\,\sigma$ confidence levels ($\Delta\chi^{2}=1$, $4$, and $9$, respectively). For comparison, the blue curves represent the NuFIT 6.0 global fit of neutrino oscillation experiments [NuFIT6] while the red curves show the improvement from the JUNO first data [JUNO1st]. In all the three panels, the improvement is quite sizable which clearly shows the importance of medium baseline reactor experiments. In all three panels, the improvement is quite sizable, which clearly demonstrates the importance of medium baseline reactor experiments. Quantitatively, from the narrowing of the $\Delta\chi^{2}$ profiles, one can estimate that JUNO has reduced the uncertainties on these effective mass limits by approximately $22.0\%$ for $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$, $22.5\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$, and $23.1\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$. The corresponding $3\,\sigma$ intervals after including the JUNO first data are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(16.36,\,21.48)\,$meV, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.268,\,1.764)\,$meV, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.484,\,3.964)\,$meV. Although JUNO does not measure the $0\nu 2\beta$ decay directly in the first several years, it plays an important role in reducing the uncertainty from particle physics side.
Translating the measurement of the half-life time $T^{0\nu}_{1/2}$ to the effective mass $\langle m_{ee}\rangle$ suffers from the uncertainty in the nuclear matrix elements (NME) $M^{0\nu}$ in Eq. (1.1). These nuclear physics uncertainties surpass those from neutrino parameter determinations. Following the framework established in [GE1], the Fig. 3 therein is updated as Fig. 3 with the current NME calculations extracted from [Agostini:2022zub]. Fig. 3 illustrates the current experimental limits on $\langle m_{ee}\rangle$ versus $T_{1/2}^{0\nu}$ for six isotopes: 48Ca, 76Ge, 82Se, 100Mo, 130Te, and 136Xe. The light blue bands indicate the nuclear uncertainties using the combined NME range of available calculations for each isotope (see Table I of [Agostini:2022zub]). Even with fixed value for the half-life time $T^{0\nu}_{1/2}$, there is no single fixed value of the effective mass $\langle m_{ee}\rangle$ to be extracted.
Fig. 3: The $0\nu 2\beta$ decay half-life time $T^{0\nu}_{1/2}$ and effective Majorana mass $|\langle m_{ee}\rangle|$ as well as their uncertainties from the nuclear and particle physics sides. Different panels are for different isotopes. The light-blue band indicates the nuclear physics uncertainty when translating the measured half-life $T^{0\nu}_{1/2}$ into a sensitivity on the $0\nu 2\beta$ effective mass $|\langle m_{ee}\rangle|$. Current experimental limits on $T^{0\nu}_{1/2}$ (90% C.L.) are shown as solid vertical lines with the region to the left of each solid line is excluded while the future projections are shown as vertical dashed lines. The horizontal bands show the effective mass limits for the vanishing lightest neutrino mass ($m_{\text{lightest}}\to 0$): $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$, with central values indicated by the red solid lines. For comparison, the orange band shows the $3\,\sigma$ ranges from the current global fit while the blue band is for the combined result of global fit plus JUNO.
The current experimental sensitivities on the half-life time $T^{0\nu}_{1/2}$ at 90% C.L. are represented by the solid vertical lines in Fig. 3. Regions to the left-hand side of the vertical line have been excluded. Among various measurements, the most stringent constraints come from KamLAND-Zen (136Xe) with $T_{1/2}^{0\nu}>2.0\times 10^{26}$ yr [KamLAND-Zen:2024eml, KamLAND-Zen:2022tow, KamLANDZen:2016pfg] and GERDA (76Ge) with $T_{1/2}^{0\nu}>1.2\times 10^{26}$ yr [GERDA:2017foj]. Other competitive bounds include MAJORANA DEMONSTRATOR (76Ge, $T_{1/2}^{0\nu}>3.9\times 10^{25}$ yr) [Majorana:2022udl], EXO-200 (136Xe, $T_{1/2}^{0\nu}>2.4\times 10^{25}$ yr) [EXO200:2019rkq], CUORE (130Te, $T_{1/2}^{0\nu}>4.8\times 10^{25}$ yr) [CUORE:2022xwt], CUPID-Mo (100Mo, $T_{1/2}^{0\nu}>9.7\times 10^{24}$ yr) [CUPID:2022okx], and CUPID-0 (82Se, $T_{1/2}^{0\nu}>4.4\times 10^{24}$ yr) [CUPID:2022okx].
Comparing these experimental limits with the horizontal bands in Fig. 3, we observe that the current experiments are beginning to probe the IO parameter space, approaching the lower edge of $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$. However, they remain several orders of magnitude away from the NO region that extends from $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ to $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$, in the limit of vanishing lightest neutrino mass ($m_{\text{lightest}}\to 0$). The orange and blue bands represent the oscillation parameter uncertainties for each of these predictions, showing how the JUNO measurement (blue band) narrows the allowed ranges compared to current global fit alone (orange band). For NO, the situation is more challenging. The maximum allowed value $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$ might be accessible to the next-generation experiments while the minimum value $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ lies at much lower masses, requiring significantly better sensitivities than currently projected.
The effort for next-generation experiments promises a qualitative leap in sensitivity, represented by the dashed vertical lines in Fig. 3, with the flagship experiments nEXO (136Xe) [nEXO1, nEXO2] and LEGEND-1000 (76Ge) [LEGEND]. Both are projected to reach an unprecedented half-life sensitivity of $T_{1/2}^{0\nu}\sim 10^{28}$ yr. As illustrated by the intersection of their dashed vertical lines with the light blue NME uncertainty band, the corresponding $\langle m_{ee}\rangle$ sensitivity ranges are $(6\sim 27)$ meV for nEXO and $(9\sim 21)$ meV for LEGEND-1000. In addition, the CUPID experiment [CUPID1, CUPID2] targeting 100Mo and 82Se aims for $T_{1/2}^{0\nu}\sim 10^{27}$ yr which can translate to $\langle m_{ee}\rangle=(12\sim 34)$ meV. We also include in Fig. 3 the projected sensitivities of DARWIN [DARWIN2], KamLAND-Zen [KamLANDZen6], NEXT-HD [NEXT2], SNO+ [SNOplus2], AMoRE-II [AMoRE], SuperNEMO [SuperNEMO1], and LEGEND-200 [LEGEND].
While these next-generation experiments represent impressive technical achievements, their projected sensitivities would only begin to probe the IO region near $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$, assuming favorable NME calculations. Definitive exclusion of IO would require reduction of the NME uncertainties or further experimental improvements beyond the current projections. For the NO case, these upcoming experiments remain far from accessing most of the allowed parameter space. Even with optimistic assumptions, nEXO and LEGEND-1000 would only marginally reach $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$, while the region near $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ would remain inaccessible. This highlights the critical importance of reducing the oscillation parameter uncertainties though future measurements. The combination of improved $0\nu 2\beta$ sensitivity and precise oscillation data with reactor neutrino experiments, particularly from JUNO, will be essential for meaningful constraints on the effective mass and the neutrino mass ordering.
4 JUNO Impact on Majorana Triangle and Majorana CP Phases
The particle physics contribution to the $0\nu 2\beta$ decay effective mass $\langle m_{ee}\rangle$ is not just the two mixing angles ($\theta_{r}$ and $\theta_{s}$) and the two mass squared differences. There are two Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ which also come from the particle physics side. Not to say the lightest neutrino mass or the absolute mass scale. However, the neutrino oscillation experiments cannot measure neither of them. While the neutrino mass scale can be probed by the beta decay spectrum experiment KATRIN and the cosmological observations as summarized in Sec. 4.1, the only chance for measuring the two Majorana CP phases comes from the $0\nu 2\beta$ decay itself. We will investigate how the improved oscillation parameter measurements affect the prospect of measuring the two Majorana CP phases through the Majorana triangle in Sec. 4.2.
4.1 Cosmological Measurement of Neutrino Mass Scale and Preference for NO
Fig. 4: (Left) Cosmological constraints on the neutrino mass sum $\sum_{i}m_{i}$ at 95% C.L. with vertical line for the exiting observations from CMB (solid) and DESI BAO (dashed). For comparison, the projected sensitivities at the future EUCLID is shown as dash-dotted line. (Right) The current KATRIN (solid) and its design (dashed) sensitivities on the beta decay effective mass $m_{\beta}$. For comparison, the expected PROJECT 8 sensitivity is also shown as a vertical dash-dotted line. The uncertainties of the effective Majorana mass $\langle m_{ee}\rangle$ from the particle physics side are also shown according to the the color scheme of Fig. 1.
The cosmological constraints on the neutrino mass sum $\sum_{i}m_{i}$ provide complementary information to laboratory searches. Especially, the DESI BAO measurements with CMB data gives the most stringent current bound, $\sum_{i}m_{i}<0.082$ eV at $95\%$ C.L. [DESI2024b, DESI2024]. For reference, the combination of CMB data from Planck, the non-DESI Baryon Acoustic Oscillations (BAO) observation, Cosmic Chronometers (CC), and Pantheon, yields $\sum_{i}m_{i}<0.12$ eV at $95\%$ C.L. [Planck:2018vyg]. These bounds are indicated in the left panel of Fig. 4 with vertical lines. The current cosmological observations have already went beyond the IO regime to probe its NO counterpart. The future cosmological surveys, such as EUCLID, are expected to reach the sensitivity of $\sum_{i}m_{i}<0.056$ eV [EUCLID2020, EUCLID2022] which would directly test the global lower limit of the neutrino mass sum.
[…]
The kinematic constraint $m_{\beta}<0.45$ eV at $90\%$ C.L. [KATRIN2022, KATRIN2024] on the beta decay effective mass $m_{\beta}\equiv\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ from the KATRIN experiment provides a direct and model-independent measurement of the neutrino mass scale. Although the beta decay constraint is not as strong as the cosmological observation, we also show them in the right panel of Fig. 4 for comparison. The current laboratory measurements probe the cosmologically allowed mass range, but still remain about an order of magnitude above the sensitivity required to test the IO scenario. The KATRIN experiment is expected to reach its design sensitivity $m_{\beta}\sim 0.2$ eV at $90\%$ C.L. after several years of data taking [KATRINFuture]. The next-generation experiments such as PROJECT 8 are projected to reach the sensitivity of $m_{\beta}\sim 0.04$ eV [Project8, Project8Roadmap], which would probe the majority of the IO regime.
Fig. 5: Allowed regions in the complex plane of the effective mass $\langle m_{ee}\rangle$ (real vs. imaginary parts) for NO. Three scenarios are presented: (Left) $m_{\text{lightest}}\in[1,50]$ meV with a flat prior and no cosmological bounds on $\sum_{i}m_{i}$; (Center) $m_{\text{lightest}}\in[1,50]$ meV with a prior from the pre-DESI CMB cosmological bounds on $\sum_{i}m_{i}$; and (Right) $m_{\text{lightest}}\in[1,50]$ meV with a prior from DESI BAO + CMB cosmological bounds on $\sum_{i}m_{i}$. The colored contours represent different confidence levels: red (68% C.L.), purple (95% C.L.), and cyan ($99\%$ C.L.). The black circle around the origin indicates the funnel region with $|\langle m_{ee}\rangle|<1$ meV.
As summarized in Sec. 2 and in particular Eq. (2.4), the effective mass $\langle m_{ee}\rangle$ can be viewed as a vector sum of three contributions in the complex plane. With our phase convention $\delta_{\rm M2}=0$, the vector $\overrightarrow{L_{2}}$ lies along the real axis while $\overrightarrow{L_{1}}$ and $\overrightarrow{L_{3}}$ rotate with the two physical Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$. This geometrical configuration is finally determined by the allowed ranges for $|U_{ei}|^{2}$ from oscillation experiments, the relative magnitudes of the triangle sides $L_{i}=m_{i}|U_{ei}|^{2}$, and the absolute mass scale $m_{\text{lightest}}$.
The behavior of $\langle m_{ee}\rangle$ in the complex plane reveals the fundamental difference between the two mass orderings. For NO, the mass hierarchy $m_{1}<m_{2}\ll m_{3}$ allows these three vectors to form a closed triangle that can reach complete cancellation, $\langle m_{ee}\rangle=0$, for specific combinations of the two Majorana CP phases [Vissani:1999tu, Xing:2014yka, GE2]. In contrast, for IO with $m_{3}\ll m_{1}<m_{2}$, the relative magnitudes of the vector sides prevent such closure, making $\langle m_{ee}\rangle=0$ geometrically impossible regardless of the Majorana phases. Such fundamental difference manifests topologically in the complex $\langle m_{ee}\rangle$ plane for NO.
Fig. 5 shows the allowed regions for the effective mass $\langle m_{ee}\rangle$ with the horizontal and vertical axes denoting its real and imaginary parts, respectively. As the Majorana phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ vary over their full range $0,2\pi]$, the allowed values of $\langle m_{ee}\rangle$ trace out a characteristic circle region, which can be understood with the geometrical picture. The real $L_{2}$ first shifts to the central point to the right. Based on this, the vector $\overrightarrow{L_{1}}$ with length $L_{1}=m_{1}c^{2}_{s}c^{2}_{r}$ rotates around the shifted epicenter and the third vector $\overrightarrow{L_{3}}$ with shorter length $L_{3}=m_{3}s^{2}_{r}$ further scanning the whole plane. Vertically, the allowed regions are symmetric around the horizontal axis which is a manifestation of the fact that the imaginary part can flip sign when the two Majorana CP phases change their signs simultaneously.
Across the three panels of Fig. 5, the size of the circles differ considerably. The major factor is the allowed range of the neutrino mass scale. For comparison, the left panel shows the results obtained with even distribution (flat prior) of the lightest mass in the range of $m_{1}\in[1,50]$ meV. Once the cosmological constraints are imposed in the center and right panels, the allowed circle regions shrink. The effect is not that significant for the case with only pre-DESI constraints since the limit $\sum_{i}m_{i}<0.12$ eV 95% C.L. [Planck:2018vyg] translates to $m_{1}\lesssim 30$ meV which is roughly the size of circle. To make the connection more direct, the corresponding upper limit for the NO effective mass in Eq. (2.8) with a nonzero $m_{1}$ is,
$\langle m_{ee}\rangle^{\rm NO}_{\rm max}(m_{1})$ $=m_{1}c^{2}_{s}c^{2}_{r}+\sqrt{\Delta m^{2}_{s}+m_{1}^{2}}c^{2}_{r}s^{2}_{s}+\sqrt{\Delta m^{2}_{a}+m_{1}^{2}}s^{2}_{r}\approx 31\,\mbox{meV}.$ (4.1)
However, once the DESI constraint $\sum m_{i}<0.082$ eV at 95% C.L.[DESI2024b, DESI2024] is imposed, the allowed circle significantly shrinks to $\langle m_{ee}\rangle^{\rm NO}_{\rm max}(m_{1})\approx 16$ meV with $m_{1}\lesssim 14$ meV.
A very important feature is that the vanishing effective mass scenario, $\langle m_{ee}\rangle=0$ which corresponds to the origin of the plots, is always allowed by the current constraints. This because the funnel or throat region shown in the left panel of Fig. 1 is covered by the sampling regions, $m_{1}\in[1,50]$ meV. Even the DESI result with $m_{1}\lesssim 14$ meV still covers this funnel region as clearly indicated by the dashed line in the left panel of Fig. 4. Most importantly, the funnel region $|\langle m_{ee}\rangle|<1$ meV indicated by the black circle around the origin is actually in the most probable range (red band). With more stringent bound from DESI, the relative size of the funnel region with respect to the whole circle actually increases. It indicates that the chance for the $0\nu 2\beta$ decay effective mass $\langle m_{ee}\rangle$ to fall into the funnel region is actually increasing.
4.2 Determination of Two Majorana CP Phases
Fig. 6: Allowed regions in the complex plane of $\langle m_{ee}\rangle$ (real vs. imaginary parts) for NO with the lightest mass $m_{1}=3$, 4, 5, 6 meV. The filled color regions represent the 68% (red), 95% (purple), and 99% (cyan) confidence levels. The black circles indicate the funnel region with $|\langle m_{ee}\rangle|<1$ meV (solid), 0.6 meV (dashed), and 0.3 meV (dash-dotted).
As shown above, the scenario of a vanishing effective mass $\langle m_{ee}\rangle$ or being bound from above has visible chance. To focus on such scenarios, Fig. 6 shows the allowed regions in the complex $\langle m_{ee}\rangle$ plane for NO with four representative values of the lightest mass, $m_{\text{lightest}}=(3$, 4, 5, 6) meV in different panels. The figure clearly illustrates the characteristic circle topology for NO that passes through the origin. For $m_{\text{lightest}}$ in the narrow window of $(3\sim 6)$ meV, the 68% C.L. region (red contours) covers the origin. In other words, the three vectors can arrange themselves into a closed triangle with complete cancellation at specific combinations of $\delta_{\rm M1}$ and $\delta_{\rm M3}$ to satisfy the geometrical closure conditions. A major difference from Fig. 5 is that now the circle is not fully filled but has a black region in the middle. This is because with fixed $m_{1}$, the three vector lengthes $L_{i}$ cannot adjust freely, especially for $L_{2}$ whose length determines the thickness of the doughnut.
If the cancellation among the three complex vectors $\overrightarrow{L_{i}}$ really happens, a Majorana triangle can form. Then it is possible to determine the two Majorana CP phases simultaneously as function of the three vector lengths [GE2],
$\cos\delta_{\rm M1}=-\frac{L_{1}^{2}+L_{2}^{2}-L_{3}^{2}}{2L_{1}L_{2}},\qquad\cos\delta_{\rm M3}=\frac{L_{1}^{2}-L_{2}^{2}-L_{3}^{2}}{2L_{2}L_{3}}.$ (4.2)
Note that only a single degree of freedom, as a combination of the two Majorana CP phases, can be determined for a nonzero effective mass. However, a vanishing $\langle m_{ee}\rangle$ requires both its real and imaginary parts to vanish. In other words, two independent conditions can appear in the vanishing limit to simultaneously constraint the two Majorana CP phases. This is a very rare example that one extra constraint can appear. Even if there is experimental uncertainty and hence only an upper bound on the effective mass can be extracted, $|\langle m_{ee}\rangle|<c$, it can transfer to both the imaginary and real parts, $|\mbox{Re}(\langle m_{ee}\rangle)|<c$ and $|\mbox{Im}(\langle m_{ee}\rangle)|<c$. The funnel region has been shown with black circles with $c=1$ meV (solid), 0.6 meV (dashed), and 0.3 meV (dash-dotted) in Fig. 6.
A paradox here is that if the $0\nu 2\beta$ decay is observed, one physical degree of freedom cannot be obversed. But if the effective mass $\langle m_{ee}\rangle$ vanishes to make the $0\nu 2\beta$ decay invisible, both physical Majorana CP phases can be simultaneously determined. The Majorana nature can be probed with other ways, such as the direct detection of cosmic neutrino backgrounds with tritium [Ringwald:2009bg, Vogel:2015vfa, Shergold:2021evs] or the atomic radiative emission of neutrino pairs [Dinh:2012qb, Fukumi:2012rn].
Fig. 7 illustrates this possibility by showing the two Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with different values of the lightest neutrino mass $m_{1}=3$, 4, 5, 6 meV. Each colored region represents one allowed solution: blue for 3 meV, orange for 4 meV, green for 5 meV, and red for 6 meV. These solutions are obtained by varying the oscillation parameters within $\Delta\chi^{2}<9$ and solving Eq. (4.2) for each parameter combination. Three uncertainty schemes are displayed: the previous uncertainties used in the Fig. 7 of [GE2] (dotted contours), current NuFIT 6.0 bounds [NuFIT6] (dashed contours), and the combination of NuFIT with the JUNO first measurement [JUNO1st] (thick solid contours). The progressive narrowing of these allowed regions, particularly evident when comparing the dotted with the solid contours, directly demonstrates how the reduced uncertainties of oscillation parameters translate into more precise predictions for the Majorana phases solutions.
Fig. 7: The two Majorana CP phases $(\delta_{\rm M1},\delta_{\rm M3})$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with $m_{1}=3$ meV (blue), 4 meV (orange), 5 meV (green), and 6 meV (red). Three uncertainty schemes of the oscillation parameters with $\Delta\chi^{2}<9$ are shown. The dotted contours represent results from the Fig. 7 of [GE2], dashed contours show the current NuFIT 6.0 bounds, while the thick solid contours incorporate the latest JUNO measurement.
The comparison in Fig. 7 clearly demonstrates how the reactor experiments can help mitigating the uncertainty in $0\nu 2\beta$ decay. The factor of $\sqrt{2}$ reduction in the $\theta_{s}$ uncertainty at JUNO narrows the allowed region of $\delta_{\rm M1}$ and $\delta_{\rm M3}$ for each value of $m_{1}$. This reduction is particularly pronounced because $\theta_{s}$ directly affects the relative magnitudes of $L_{1}$ and $L_{2}$ through the prefactors $c_{s}^{2}$ and $s_{s}^{2}$. The narrower bands in Fig. 7 translate directly into more precise predictions. If future $0\nu\beta\beta$ experiments do not observe a signal despite reaching sensitivities $|\langle m_{ee}\rangle|\lesssim 1$ meV, JUNO’s improved oscillation parameters would enable a much sharper determination of the Majorana CP phases under the assumption of complete cancellation.
5 Conclusion
In this paper, we have investigated the impact of the JUNO first data release on the $0\nu 2\beta$ decay effective mass and the determination of Majorana CP phases. The improved precision on the oscillation parameters, particularly almost a factor of $25.9\%$ reduction in the uncertainty of $\sin^{2}\theta_{s}$ and $37.3\%$ reduction in $\Delta m^{2}_{s}$ at $3\,\sigma$ range, directly translates into 22% (23%) reduction in the uncertainty of the minimum effective Majorana mass for IO (NO). At the $3\sigma$ confidence level, we find $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(16.36\sim 21.48)$ meV, with the narrowed range providing a more precise target for the next-generation $0\nu\beta\beta$ experiments. This improvement is crucial for the experimental design of facilities such as nEXO, LEGEND-1000, and CUPID, which aim to probe the IO parameter space. The sharpened lower bound means that a null result from these experiments, combined with JUNO’s precision measurements, would provide increasingly strong evidence against the IO scenario. Nevertheless, the JUNO experiment can provide direct measurement of the neutrino mass ordering.
Beyond the bounds on $\langle m_{ee}\rangle$, the improved measurement of oscillation parameters at JUNO can also significantly enhance our ability to determine the Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ in the scenario of complete cancellation ($\langle m_{ee}\rangle=0$) for NO. For the narrow mass window $m_{1}=3\sim 6$ meV where a vanishing effective mass can exist, the allowed regions in the $(\delta_{\rm M1},\delta_{\rm M3})$ plane are substantially reduced from to previous analyses. Although simultaneous determination of the two Majorana CP phases is possible if future $0\nu 2\beta$ experiments do not observe a signal when reaching the sensitivity of $|\langle m_{ee}\rangle|\lesssim 1$ meV, it still relies on an experiment such as JUNO to reduce uncertainties from the particle physics side.
Looking ahead, the JUNO full dataset with several years of operation is expected to achieve sub-percent precision on $\sin^{2}\theta_{s}$, which would further narrow the theoretical predictions for $\langle m_{ee}\rangle$ in both mass orderings. This improvement will be especially valuable when combined with the next-generation $0\nu 2\beta$ experiments as well as the absolute neutrino mass measurement from cosmological observations (such as DESI, EUCLID, and future CMB experiments) and beta decay experiments (such as PROJECT 8). The synergy of these three complementary measurements — the precision neutrino oscillation, the $0\nu 2\beta$ decay, and the absolute mass scale determinations — forms an essential triad for fully characterizing the nature of neutrino masses. The results presented in this work demonstrate that the precision era of neutrino physics has truly begun, with JUNO playing a central role in shaping the landscape of $0\nu 2\beta$ decay phenomenology for the coming decade.
Acknowledgements
The authors would like to thank Yue Meng for useful discussions. The authors are supported by the National Natural Science Foundation of China (12425506, 12375101, 12090060, and 12090064) and the SJTU Double First Class start-up fund (WF220442604). CFK is supported by IBS under the project code IBS-R018-D1. SFG is also an affiliate member of Kavli IPMU, University of Tokyo.